# Towards Evaluating Large Language Models on Sarcasm Understanding

### v1

@@ Abstract
In the era of large language models (LLMs), the task of “System I” - the fast, unconscious, and intuitive tasks, e.g., sentiment analysis, text classification, etc., have been argued to be successfully solved. However, sarcasm, as a subtle linguistic phenomenon, often employs rhetorical devices like hyperbole and figuration to convey true sentiments and intentions, involving a higher level of abstraction than sentiment analysis. There is growing concern that the argument about LLMs’ success may not be fully tenable when considering sarcasm understanding. To address this question, we select eleven SOTA LLMs and eight SOTA pre-trained language models (PLMs) and present comprehensive evaluations on six widely used benchmark datasets through different prompting approaches, i.e., zero-shot input/output (IO) prompting, few-shot IO prompting, chain of thought (CoT) prompting. Our results highlight three key findings: (1) current LLMs underperform supervised PLMs based sarcasm detection baselines across six sarcasm benchmarks. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. (2) GPT-4 consistently and significantly outperforms other LLMs across various prompting methods, with an average improvement of 14.0$\uparrow$. Claude 3 and ChatGPT demonstrate the next best performance after GPT-4. (3) Few-shot IO prompting method outperforms the other two methods: zero-shot IO and few-shot CoT. The reason is that sarcasm detection, being a holistic, intuitive, and non-rational cognitive process, is argued not to adhere to step-by-step logical reasoning, making CoT less effective in understanding sarcasm compared to its effectiveness in mathematical reasoning tasks.
@@
Large language models , sarcasm detection , CoT prompting
@@ 1 Introduction
Recent large language models (LLMs) have demonstrated outstanding instruction-following and in context learning abilities across various natural language processing (NLP) tasks, such as question answering [1], sentiment analysis [2], text classification [3], etc. In the era of LLMs, it has been argued that “System I” tasks - the fast, unconscious, and intuitive tasks have been successfully solved [4]. Persistent focus and efforts from both the academic and industrial sectors have primarily concentrated on what is known as “System II” tasks. Such tasks demand slow, deliberate, and multiple-step cognitive processes, including logical, mathematical, and commonsense reasoning [5].
@@
The experimental results highlight three key findings: (1) Current LLMs underperform supervised PLMs based sarcasm detection baselines across six sarcasm benchmarks. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. (2) GPT-4 consistently and significantly outperforms other LLMs across various prompting methods, with an average improvement of 14.0$\uparrow$. Claude 3 and ChatGPT demonstrate the next best performance after GPT-4. (3) Few-shot IO prompting method outperforms the other two methods: zero-shot IO and few-shot CoT, with an average improvement of 4.5$\uparrow$. The reason is that sarcasm detection, being a holistic, intuitive, and non-rational cognitive process, is argued not to adhere to step-by-step logical reasoning, making CoT less effective in understanding sarcasm compared to its effectiveness in mathematical reasoning tasks. These findings underscore the complex nature of sarcasm detection and the current limitations of LLMs, while also highlighting the further need of stronger LLMs.
@@
  • Our work is the first to investigate the performance of LLMs in understanding sarcasm.
@@ 2 Related Work
This section reviews two lines of research that form the basis of this work: CoT prompting and sarcasm detection.
@@ 2.1 Sarcasm Detection
Sarcasm detection is habitually treated as a text classification task, where the target is to identify whether the given text is sarcastic or not [16]. It has evolved from early rule based and statistical learning based approaches to traditional neural methods, such as CNN, RNN, and further advanced to modern neural methods epitomized by Transformer models. In early stage, the rule based approaches infer the overall sarcasm polarity based on the refined sarcasm rules, such as the occurrence of the interjection word [9]. Statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, RF, NB, etc., to extract patterns and relationships within the data [17].
As deep learning based architectures have shown the superiority over statistical learning, numerous base neural networks, e.g., such as CNN [18], LSTM [19], GCN [20], etc., have been predominantly utilized during the middle stage of sarcasm detection research, aiming to learn and extract complex features in an end-to-end fashion. As the field of deep learning continues to evolve, sarcasm detection research has stepped into the era of pre-trained language models (PLMs). An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations. For example, Liu et al. proposed a dual-channel framework by modeling both literal and implied sentiments separately [21]. They also constructed two conflict prompts to elicit PLMs to generate the sarcasm polarity [22]. Qiao et al. presented a mutual-enhanced incongruity learning network to take advantage of the underlying consistency between the two modules to boost the performance [23]. Tian et al. proposed a dynamic routing Transformer network to activate different routing transformer modules for modeling the dynamic mechanism in sarcasm detection [24].
However, the above-mentioned works still focus on how to utilize PLMs to extract effective features, without leveraging the extraodinary context learning capabilities of LLMs. In contrast, this paper employs three prompting methods to make the first attempt to explore the potential of prompting LLMs in sarcasm detection.
@@ 2.2 Large Language Models
LLMs are advanced language models characterized by their substantial parameter sizes and exceptional learning capabilities  [25]. In recent years, the community of natural language processing (NLP) has witnessed substantial progress largely due to the development of large language models (LLMs). These models excel in areas such as in-context learning, few-shot prompting, and following complex instructions. OpenAI has been at the forefront of this innovation, notably with the introduction of transformative models such as ChatGPT and GPT-4. Nevertheless, the exclusive nature of these technologies has led to the emergence of various LLM versions, which often incorporate tens or even hundreds of billions of parameters [26]. We categorize these LLMs into two groups based on their specialization: general LLMs and specialized LLMs.
General LLMs are designed for versatility across a wide spectrum of NLP tasks. Prominent examples of these models are GPT-4, ChatGLM 4, LLaMA $3$https://llama.meta.com/, PanGu-$\Sigma$ [27], Baichuan 3444https://www.baichuan-ai.com, etc. Such LLMs often perform well across a range of tasks, but their potentials in specific domains await further explore. In contrast, specialized LLMs are fine-tuned for specific tasks via task-specific architectures and knowledge, allowing them to achieve higher performance. For example, Zhang et al. proposed a fine-tuned context and emotion knowledge tuned LLM for emotion recognition in conversations [2]. They also presented RGPT, an adaptive boosting framework tailored to produce a specialized text classification LLM by recurrently ensembling a pool of strong base learners [3, 28]. Moreover, an increasing number of studies are raging over medical, law, finance and education domains, e.g., HuaTuo [29], FinGPT [30], ChatLaw [31], etc.
@@ 3.1 Task Definition
Sarcasm detection is transformed as a conditional generative task, where the output $\mathcal{Y}$ will be the labels. Given a set of input texts $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N}\}$ where each document $x_{i}$ is augmented with a designed prompt ${Prompt}_{i}\in\mathcal{P}$ that provides contextual guidance, i.e., ${Prompt}_{i}={INS}_{i}\oplus x_{i}$, where ${INS}_{i}$ represents the task instruction, $\mathcal{P}$ represents the prompt set. Our task is to let a LLM $\mathcal{M}$ map an input document to its target label: $\mathcal{M}(\mathcal{X},\mathcal{P},\theta)\rightarrow\mathcal{Y}$, where $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{N}\}$ denotes the label sequence generated by the LLM $\mathcal{M}$ We formulate the problem as:
@@
Figure 2: The proposed evaluation framework.
@@ 3.2 Evaluation Approach
Fig. 2 shows the overview of our proposed evaluation framework. Initially, we take both sarcastic and non-sarcastic samples as inputs. Subsequently, we design prompts tailored specifically for sarcasm. This involves constructing task instructions and providing contextual demonstrations. Our evaluation utilizes three distinct prompting approaches: zero-shot IO prompting, few-shot IO prompting, and CoT prompting. These methods are implemented to assess how effectively LLMs generate responses that are both valid and rigorously aligned with the subtleties of sarcastic expressions. Finally, we analyze and compare the outputs generated by the LLMs against the ground truth to evaluate the models’ proficiency in understanding and detecting sarcasm.
@@ 3.3 Prompt Construction
(3) Demonstration. It provides specific examples of how to analyze text for sarcasm, showing both sarcastic and non-sarcastic responses to similar situations. It will guide the LLMs in recognizing subtle cues and patterns that identiy sarcasm. In addition, it also provides an output format that LLM’s outputs should to follow. Note that demonstrations are only needed for the few-shot setup, but not for the zero-shot setup.
@@ 3.4 Three Prompting Approaches
Zero-shot Input-Output Prompting. Zero-shot IO prompting refers to the scenario where a language model (LM) generates the output $y$ based purely on the input $x$, without any additional examples or guidance. Mathematically, this can be represented as:
@@
Few-shot Input-Output Prompting. Few-shot IO prompting involves providing the language model with a limited number of examples or demonstrations of the task before predicting the output for a new input $x$. This can be formalized as:
@@
Chain-of-Thought Prompting. Few-shot CoT prompting extends the idea of few-shot prompting by explicitly modeling intermediate reasoning steps between the input $x$ and the final output $y$. The language model is guided through a sequence of logical steps $z_{1},z_{2},\dots,z_{n}$, leading to the final answer. This process is described as:
@@ 4 Experiments
4.1 Experiment Setups
Datasets. Six benchmarking datasets are selected as the experimental beds which encompass the most significant datasets used in sarcasm detection tasks. They are IAC-V1 [33], IAC-V2 [34], Ghosh [35], iSarcasmEval [36], Riloff [37] and SemEval 2018 Task 3 [38].
@@ 4.1 Experiment Setups
Ghosh consists of 51,189 tweets (24,453 sarcastic tweets and 26,736 non-sarcastic tweets) in which sarcastic tweets are automatically collected from Twitter using user’s self-declaration of sarcasm/irony with sarcastic and ironic hashtags (e.g. #irony, #sarcasm). In this work, we have conducted a thorough double-check of this dataset and successfully filtered out 7,804 noisy tweets.
@@
SemEval 2018 Task 3 is collected using irony-related hashtags (i.e. #irony, #sarcasm, #not) and are subsequently manually annotated to minimise the amount of noise in the corpuses. It emphasize the challenges inherent in identifying sarcasm within the constraints of MUStARD’s concise format, and highlight the importance of context and linguistic subtleties in recognizing sarcasm. The statistics for each dataset are shown in Table 1.
Evaluation metrics. To make a fair comparison, we employ precision (P), recall (R), accuracy (Acc) and F1 as evaluation metrics. For each method, we run five random seeds and report the average result of the test sets. The metrics are written as:
$P=\frac{TP}{TP+FP}$ (2)
$R=\frac{TP}{TP+FN}$
$F1=\frac{2\ast P\ast R}{P+R}$
$Acc=\frac{TP+TN}{TP+FN+FP+TN}$
@@
Implementation details. We use the official implementations and hyper-parameters for all DLMs and PLMs based approaches. For example, all weight matrices are given their initial values by sampling from a uniform distribution $U(-0.1,0.1)$. The learning rate is set to 5e-5. The batch size is set to 32 and the number of epochs is set to 30. The dropout rate is set to 0.5. In constrast, the LLMs based methods are implemented with the respective official Python API library and the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers. All the experiments are conducted on an NVIDIA GeForce GTX 4090 GPU.
@@ 4.2 Compared Baselines
A wide range of strong baselines are included for comparison including traditional deep learning models (DLMs), pre-trained language models (PLMs) and LLMs. They are:
@@
    (3) LSTM takes word embeddings as input and take the hidden representation of each word for sarcasm classification.
    (4) bi-LSTM [41] implements an bidirectional LSTM to learn both forward and backward long-range dependency information.
@@
Prompting approaches. Three prompting approaches, i.e., zero-shot IO, few-shot IO and CoT are adopted to elicit LLMs to produce correct answers, as shown in Fig. 3.
@@ 4.3 Results and Analysis of Zero-Shot IO Prompting
(1) In the zero-shot IO setting, PLMs demonstrate outstanding performance, significantly surpassing traditional deep learning methods. For instance, RoBERTa achieves an average F1 score of 71.3 across all datasets, while DC-Net-RoBERTa reaches 71.5. In contrast, traditional models like Bi-LSTM and AT-LSTM score lower, with F1 scores of 66.7 and 65.3, respectively. PLMs outperform traditional methods by approximately 7.8$\uparrow$. This superiority stems from the extensive pretraining that PLMs undergo on large-scale data, enabling them to better capture contextual information and adapt to various tasks. Although traditional deep learning models perform adequately on specific tasks, their lack of large-scale pretraining support limits their generalization capabilities, resulting in weaker performance.
(2) Despite that current Large Language Models have not yet fully outperformed PLM-based supervised learning methods, their potential is evident in specific tasks. For example, GPT-4 Turbo excels in the IAC-V1 and SemEval Task 3 datasets, achieving F1 scores of 78.7 and 76.5, respectively, significantly outperforming RoBERTa’s scores of 69.9 and 72.0, with the margin of 8.8% and 4.5$\uparrow$. However, LLMs fall short in datasets like iSarcasmEval and Riloff, where their performance does not meet expectations. The average F1 score shows that GPT-4 Turbo’s difference from RoBERTa in the IAC-V2 and Ghosh datasets is relatively small. This performance variation may be due to LLMs’ advantage in handling long texts and complex semantics, but their generalization capabilities are weaker when facing domain-specific or fine-grained tasks without specialized tuning. Therefore, while LLMs show strong potential in some tasks, their consistency and generalization across different tasks still need improvement.
(3) Among all 11 large language models, GPT-4 Turbo demonstrates a comprehensive lead. It performs exceptionally well across multiple datasets, particularly in IAC-V1 and SemEval Task 3, where it achieves F1 scores of 78.7 and 76.5, significantly higher than Claude-3-haiku’s 77.0 and 57.9. Additionally, in the Ghosh dataset, GPT-4 Turbo reaches an F1 score of 82.2, far surpassing other models like ChatGPT (71.4), Claude-3-haiku (71.8) and LLaMA 3-8B (66.8). A significant reason for GPT-4’s superiority over other models is its larger parameter size, which enhances its human language comprehension and reasoning capabilities. In contrast, most other models are smaller, with around 7B/8B parameters, which limits their ability to handle complex tasks. Moreover, GPT-4’s ability to adapt to various data types and text structures across different tasks further contributes to its outstanding performance. Thus, GPT-4’s strong results underscore its leading position among LLMs and its powerful generalization capabilities.
@@
4.4 Few-Shot v/s Zero-Shot
Since the above experiments are mainly based on a zero-shot IO setting, we are curious of whether the conclusions also apply in a few-shot scenario. Therefore, we perform few-shot experiments to evaluate whether the LLMs can perform better when a limited number of contextual examples are available. We show the main results in Table 3 and Fig. 5, we sample $k=2$ examples (including one sarcastic example and one non-sarcastic exmaple) from the training set via KNN search.
@@ 4.4 Few-Shot v/s Zero-Shot
Figure 5: The comparison between zero-shot IO prompting and few-shot IO prompting.
Figure 6: The comparison between few-shot IO prompting and few-shot CoT prompting.
@@
(2) A series of LLMs exhibit slight performance improvements compared to the zero-shot setting. For instance, ChatGPT’s F1 score on the IAC-V1 dataset rises from 70.0 in zero-shot to 73.2 in few-shot, reflecting a 4.6% improvement. Similarly, ChatGLM 3-6B shows an increase in its F1 score on IAC-V1 from 52.0 in zero-shot to 52.7 in few-shot, representing a 1.3% gain. Mistral-7B displays a more notable improvement on the iSarcasmEval dataset, where its F1 score advances from 26.7 in zero-shot to 28.0 in few-shot, indicating a 2.9% enhancement. Even though the extent of enhancement differs across models and datasets, the few-shot setup generally contributes to better outcomes.

### v2

# SarcasmBench: Towards Evaluating Large Language Models on Sarcasm Understanding

@@ Abstract
In the era of large language models (LLMs), the task of “System I” - the fast, unconscious, and intuitive tasks, e.g., sentiment analysis, text classification, etc., have been argued to be successfully solved. However, sarcasm, as a subtle linguistic phenomenon, often employs rhetorical devices like hyperbole and figuration to convey true sentiments and intentions, involving a higher level of abstraction than sentiment analysis. There is growing concern that the argument about LLMs’ success may not be fully tenable when considering sarcasm understanding. To address this question, we select eleven SOTA LLMs and eight SOTA pre-trained language models (PLMs) and present comprehensive evaluations on six widely used benchmark datasets through different prompting approaches, i.e., zero-shot input/output (IO) prompting, few-shot IO prompting, chain of thought (CoT) prompting. We build SarcasmBench, the first benchmark to comprehensively evaluate LLMs on sarcasm understanding. Our results highlight three key findings: (1) current LLMs underperform supervised PLMs based sarcasm detection baselines. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. (2) GPT-4 consistently and significantly outperforms other LLMs across various prompting methods, with an average improvement of 14.0$\uparrow$. (3) Few-shot IO prompting method outperforms the other two methods: zero-shot IO and few-shot CoT. We hope this paper can provide some guidance for subsequent researchers to better solve sarcasm detection in the era of LLMs.
@@
Large language models , sarcasm detection , prompting strategy
journal: INFORMATION PROCESSING & MANAGEMENT
@@ 1 Introduction
Recent large language models (LLMs) have demonstrated outstanding instruction-following and in-context learning abilities across various natural language processing (NLP) tasks, such as question answering [1], sentiment analysis [2], text classification [3], etc. In the era of LLMs, it has been argued that “System I” tasks - the fast, unconscious, and intuitive tasks have been successfully solved [4]. Persistent focus and efforts from both the academic and industrial sectors have primarily concentrated on what is known as “System II” tasks. Such tasks demand slow, deliberate, and multiple-step cognitive processes, including logical, mathematical, and commonsense reasoning [5].
@@
The experimental results highlight three key findings: (1) Current prompting-based LLMs underperform supervised PLMs across six benchmark datasets. This suggests that significant efforts are still required to improve LLMs’ understanding of human sarcasm. (2) GPT-4 consistently and significantly outperforms other LLMs across various prompting methods, with an average improvement of 14.0$\uparrow$. Claude 3 and ChatGPT demonstrate the next best performance after GPT-4. (3) Few-shot IO prompting method outperforms the other two methods: zero-shot IO and few-shot CoT, with an average improvement of 4.5$\uparrow$. The reason is that sarcasm detection, being a holistic, intuitive, and non-rational cognitive process, is argued not to adhere to step-by-step logical reasoning, making CoT less effective in understanding sarcasm compared to its effectiveness in mathematical reasoning tasks. This also aligns closely with the key findings of the SarcasmCue framework [16]. These findings underscore the complex nature of sarcasm detection and the current limitations of LLMs, while also highlighting the further need of stronger LLMs.
@@
  • This is the first work that comprehensively evaluates the performance of LLMs in understanding sarcasm.
@@ 2 Related Work
This section reviews two lines of research that form the basis of this work: sarcasm detection and large language models.
@@ 2.1 Sarcasm Detection
Sarcasm detection aims to identify whether the given text is sarcastic or not [17]. It has evolved from early rule based and statistical learning based approaches to traditional neural methods, such as CNN, RNN, and further advanced to modern neural methods epitomized by Transformer models. In the early stage, the rule based approaches infer the overall sarcasm polarity based on the refined sarcasm rules, such as the occurrence of the interjection word [9]. Statistical learning based approaches mainly employ statistical learning techniques, e.g., SVM, RF, NB, etc., to extract patterns and relationships within the data [18].
As deep learning has shown the superiority over statistical learning, numerous base neural networks, e.g., such as CNN [19], LSTM [20], GCN [21], etc., have been predominantly utilized during the middle stage of sarcasm detection research, aiming to learn and extract complex features in an end-to-end fashion. As the field of deep learning continues to evolve, sarcasm detection research has stepped into the era of pre-trained language models (PLMs). An increasing number of researchers are designing sophisticated PLM architectures to serve as encoders for obtaining effective text representations. For example, Liu et al. [22] proposed a dual-channel framework by modeling both literal and implied sentiments separately [22]. They also constructed two conflict prompts to elicit PLMs to generate the sarcasm polarity [23]. Qiao et al. presented a mutual-enhanced incongruity learning network to take advantage of the underlying consistency between the two modules to boost the performance [24]. Tian et al. proposed a dynamic routing Transformer network to activate different routing transformer modules for modeling the dynamic mechanism in sarcasm detection [25].
However, the above-mentioned works still focus on how to utilize PLMs to extract effective features, without leveraging the extraordinary understanding capabilities of LLMs. In contrast, this paper employs three prompting methods to make the first attempt to explore the potential of prompting LLMs in sarcasm detection.
@@ 2.2 Large Language Models
LLMs are advanced language models characterized by their substantial parameter sizes and exceptional learning capabilities  [26]. In recent years, the community of natural language processing (NLP) has witnessed substantial progress largely due to the development of LLMs. These models excel in areas such as in-context learning, few-shot prompting, and following complex instructions. OpenAI has been at the forefront of this innovation, notably with the introduction of transformative models such as ChatGPT and GPT-4. Nevertheless, the exclusive nature of these technologies has led to the emergence of various LLM versions, which often incorporate tens or even hundreds of billions of parameters [27]. We categorize these LLMs into two groups based on their specialization: general LLMs and specialized LLMs.
General LLMs are designed for versatility across a wide spectrum of NLP tasks. Prominent examples of these models are GPT-4, ChatGLM 4, LLaMA $3$https://llama.meta.com/, PanGu-$\Sigma$ [28], Baichuan 3444https://www.baichuan-ai.com, etc. Such LLMs often perform well across a range of tasks, but their potentials in specific domains await further exploration. In contrast, specialized LLMs are fine-tuned for specific tasks via task-specific architectures and knowledge, allowing them to achieve higher performance. For example, Zhang et al. proposed a fine-tuned context and emotion knowledge tuned LLM for emotion recognition in conversations [2]. They also presented RGPT, an adaptive boosting framework tailored to produce a specialized text classification LLM by recurrently ensembling a pool of strong base learners [3, 29].
@@ 3.1 Task Definition
Sarcasm detection is transformed as a conditional generative task, where the output $\mathcal{Y}$ will be the labels. Given a set of input texts $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N}\}$ where each document $x_{i}$ is augmented with a designed prompt ${Prompt}_{i}\in\mathcal{P}$ that provides contextual guidance, i.e., ${Prompt}_{i}={INS}_{i}\oplus x_{i}$, where ${INS}_{i}$ represents the task instruction, $\mathcal{P}$ represents the prompt set. Our task is to let a LLM $\mathcal{M}$ map an input document to its target label: $\mathcal{M}(\mathcal{X},\mathcal{P},\theta)\rightarrow\mathcal{Y}$, where $\mathcal{Y}=\{y_{1},y_{2},\ldots,y_{N}\}$ denotes the label sequence generated by the LLM $\mathcal{M}$. We formulate the problem as:
@@
Figure 2: The proposed evaluation framework for LLMs and PLMs/DLMs.
@@ 3.2 Evaluation Approach
Fig. 2 shows the overview of our proposed evaluation framework. Initially, we take both sarcastic and non-sarcastic samples as inputs. Subsequently, we design prompts tailored specifically for sarcasm. This involves constructing task instructions and providing examples for few-shot learning. Our evaluation utilizes three distinct prompting approaches: zero-shot IO prompting, few-shot IO prompting, and CoT prompting. These methods are implemented to assess how effectively LLMs generate responses that are both valid and rigorously aligned with the subtleties of sarcastic expressions. Finally, we analyze and compare the outputs generated by the LLMs against the ground truth to evaluate the models’ proficiency in understanding and detecting sarcasm.
@@ 3.3 Prompt Construction
(3) Demonstration. It provides specific examples of how to analyze text for sarcasm, showing both sarcastic and non-sarcastic responses to similar situations. It will guide the LLMs in recognizing subtle cues and patterns that identify sarcasm. In addition, it also provides an output format that LLM’s outputs should follow. Note that demonstrations are only needed for the few-shot setup, but not for the zero-shot setup.
@@ 3.4 Three Prompting Approaches
Zero-shot IO Prompting refers to the scenario where LLMs generate the output $y$ based purely on the input $x$, without any additional examples or guidance. Mathematically, this can be represented as:
@@
Few-shot IO Prompting involves providing LLMs with a limited number of examples or demonstrations of the task before predicting the output for a new input $x$. This can be formalized as:
@@
Few-shot CoT prompting extends the idea of few-shot prompting by explicitly modeling intermediate reasoning steps between the input $x$ and the final output $y$. The LLM is guided through a sequence of logical steps $z_{1},z_{2},\dots,z_{n}$, leading to the final answer. This process is described as:
@@ 4 Experiments
4.1 Dataset
Datasets. Six benchmark datasets are selected as the experimental beds which encompass the most significant datasets used in sarcasm detection tasks. They are IAC-V1 [31], IAC-V2 [32], Ghosh [33], iSarcasmEval [34], Riloff [35] and SemEval 2018 Task 3 [36].
@@ 4.1 Dataset
Ghosh consists of 51,189 tweets (24,453 sarcastic tweets and 26,736 non-sarcastic tweets) in which sarcastic tweets are automatically collected from Twitter using the user’s self-declaration of sarcasm/irony with sarcastic and ironic hashtags (e.g. #irony, #sarcasm). In this work, we have conducted a thorough double-check of this dataset and successfully filtered out 7,804 noisy tweets.
@@
SemEval 2018 Task 3 is collected using irony-related hashtags (i.e. #irony, #sarcasm, #not) and are subsequently manually annotated to minimize the amount of noise in the corpora. It emhpasizes the challenges inherent in identifying sarcasm within the constraints of MUStARD’s concise format and highlights the importance of context and linguistic subtleties in recognizing sarcasm. The statistics for each dataset are shown in Table 1.
4.2 Evaluation Metrics
To make a fair comparison, we employ precision (P), recall (R), accuracy (Acc) and F1 as evaluation metrics. For each method, we run five random seeds and report the average result of the test sets. The metrics are written as:
$P=\frac{TP}{TP+FP},$ (2)
$R=\frac{TP}{TP+FN},$
$F1=\frac{2\ast P\ast R}{P+R},$
$Acc=\frac{TP+TN}{TP+FN+FP+TN},$
@@ 4.2 Evaluation Metrics
4.3 Implementation Details
We use the official implementations and hyper-parameters for all traditional deep learning models (DLMs) and PLMs based approaches. For example, all weight matrices are given their initial values by sampling from a uniform distribution $U(-0.1,0.1)$. The learning rate is set to 5e-5. The batch size is set to 32 and the number of epochs is set to 30. The dropout rate is set to 0.5. In contrast, the LLMs based methods are implemented with the respective official Python API library and the Hugging Face Transformers library [^5] https://huggingface.co/docs/transformers. All the experiments are conducted on an NVIDIA GeForce GTX 4090 GPU.
@@ 4.4 Compared Baselines
A wide range of strong baselines are included for comparison including DLMs, PLMs and LLMs. They are:
@@
    (3) LSTM takes word embeddings as input and takes the hidden representation of each word for sarcasm classification.
    (4) Bi-LSTM [39] implements a bidirectional LSTM to learn both forward and backward long-range dependency information.
@@ 4.5 Results and Analysis of Zero-Shot IO Prompting
(1) In the zero-shot IO setting, PLMs demonstrate outstanding performance, significantly surpassing traditional DLMs. For instance, RoBERTa achieves an average F1 score of 71.3 across all datasets, while DC-Net-RoBERTa reaches 71.5. In contrast, traditional models like Bi-LSTM and AT-LSTM score lower, with F1 scores of 66.7 and 65.3, respectively. PLMs outperform traditional methods by approximately 7.8$\uparrow$. This superiority stems from the extensive pretraining that PLMs undergo on large-scale data, enabling them to better capture contextual information and adapt to various tasks. Although traditional deep learning models perform adequately on specific tasks, their lack of large-scale pretraining support limits their generalization capabilities, resulting in weaker performance.
(2) Despite that current LLMs have not yet fully outperformed PLMs, their potential is evident in particular tasks. For example, GPT-4 Turbo excels in the IAC-V1 and SemEval Task 3 datasets, achieving F1 scores of 78.7 and 76.5, respectively, significantly outperforming RoBERTa’s scores of 69.9 and 72.0, with the margin of 8.8% and 4.5$\uparrow$. However, LLMs fall short in datasets like iSarcasmEval and Riloff, where their performance does not meet expectations. The average F1 score shows that GPT-4 Turbo’s difference from RoBERTa in the IAC-V2 and Ghosh datasets is relatively small. This performance variation may be due to LLMs’ advantage in handling long texts and complex semantics, but their generalization capabilities are weaker when facing domain-specific or fine-grained tasks without specialized tuning. Therefore, while LLMs show strong potential in some tasks, their consistency and generalization across different tasks still need improvement.
(3) Among all 11 LLMs, GPT-4 Turbo demonstrates a comprehensive lead. It performs exceptionally well across multiple datasets, particularly in IAC-V1 and SemEval Task 3, where it achieves F1 scores of 78.7 and 76.5, significantly higher than Claude-3-haiku’s 77.0 and 57.9. Additionally, in the Ghosh dataset, GPT-4 Turbo reaches an F1 score of 82.2, far surpassing other models like ChatGPT (71.4), Claude-3-haiku (71.8) and LLaMA 3-8B (66.8). A significant reason for GPT-4’s superiority over other models is its larger parameter size, which enhances its human language comprehension and reasoning capabilities. In contrast, most other models are smaller, with around 7B/8B parameters, which limits their ability to handle complex tasks. Moreover, GPT-4’s ability to adapt to various data types and text structures across different tasks further contributes to its outstanding performance. Thus, GPT-4’s strong results underscore its leading position among LLMs and its powerful generalization capabilities.
@@
4.6 Few-Shot vs. Zero-Shot
Since the above experiments are mainly based on a zero-shot IO setting, we are curious of whether the conclusions also apply in a few-shot scenario. Therefore, we perform few-shot experiments to evaluate whether the LLMs can perform better with a limited number of examples. We show the main results in Table 3 and Fig. 5, we sample $k=2$ examples (including one sarcastic example and one non-sarcastic example) from the training set via KNN search.
@@
Figure 5: The comparison between zero-shot IO prompting, few-shot IO prompting and CoT prompting.
@@
(2) A series of LLMs exhibit slight performance improvements compared to the zero-shot setting. For instance, ChatGPT’s F1 score on the IAC-V1 dataset rises from 70.0 in zero-shot to 73.2 in few-shot, reflecting a 4.6% improvement. Similarly, ChatGLM 3-6B shows an increase in its F1 score on IAC-V1 from 52.0 in zero-shot to 52.7 in few-shot, representing a 1.3% gain. Mistral-7B displays a more notable improvement on the iSarcasmEval dataset, where its F1 score advances from 26.7 in zero-shot to 28.0 in few-shot, indicating a 2.9% enhancement. Even though the extent of enhancement differs across models and datasets, the few-shot setup generally contributes to better outcomes. We can also see that 7B/6B LLMs has relatively based in-context learning ability and they cannot address too complex prompts. But for some business LLMs such as GPT-4, with larger model size, they have better in-context learning ability.